1 L D5 1/2NS was started to infuse at 125 mL/hr at 7:00 am. At 1300 hrs, how much volume should be remaining in the IV bag? ______mL. At this point how many grams of Dextrose have you given? _____grams. How much NS has infused?________Once the entire amount has infused, the patient will have received NS _______mg and Dextrose _______ mg. What time should the next IV bag be ready to hang?

show all work
Answer #1

Ans) 1 L D5 1/2NS D5 1/2NS 5% dextrose in half amount of normal saline (0.45% w/v of NaCl).

was started to infuse at 125 mL/hr at 7:00 am. At 1300 hrs, how much volume should be remaining in the IV bag?

250 mL.

- 7-13 means 6 hrs= 125 × 6 = 750 ml, so 1000-750 = 250 ml

At this point how many grams of Dextrose have you given? _____grams. How much NS has infused? 750 ml

On the entire amount has infused, the patient will have received NS _______mg and Dextrose _______ mg.

Half-normal saline (0.45% NaCl), often with "D5" (5% dextrose), contains 77 mEq/L of Na and Cl and 50 g/L dextrose. Quarter-normal saline (0.22% NaCl) has 39 mEq/L of Na and Cl and almost always contains 5% dextrose for osmolality reasons.

What time should the next IV bag be ready to hang? 1500 means at 3pm as 250 ml will be remaining at 1 pm so 125+125 = 250 ml in next 2 hrs means 3pm, the Infusion will end and next will have to start.